OpenStudy (anonymous):

( x + 1.2 ) + ( 3.75 +2x ) + (1.05 + 2x ) = x + 106

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

what are you trying to do, solve for x? if so you simply combine like terms. (1.2+3.75+1.05) + (x+2x+2x)= x + 106 (6)+5x=x+106 -6 -6 _________________ 5x =x+100 then subtract the x 5x= x+100 -x -x _______________ 4x=100 divide by 4 to get x by itself and you get x= 25

OpenStudy (anonymous):

alright, thank you!